### v1

@@ 4.1 Building Quantum Embeddings
Classically, one of the most well-known recent techniques for building word embeddings is word2vec (Mikolov et al., 2013a). Taking inspiration from this line of work, we propose a quantum computing implementation of word2vec.
Word2vec is a group of word embedding methods that use shallow neural networks to capture the semantic properties of words. Word2vec includes two popular methods: Continuous Bag-of-Words (CBOW) and Skip-gram. The key idea behind both methods is that words that appear in similar contexts tend to have similar meanings, and thus similar embeddings. The similarity between two words is measured by the inner product of their corresponding vectors.
CBOW and Skip-gram have different ways of learning the word embeddings. CBOW takes the context words around the target word as input and tries to predict the target word. Skip-gram does the opposite: it takes the target word as input and tries to predict the context words. The output of both methods is a probability distribution over the entire words in the vocabulary, which is computed using the softmax function. This can be computationally expensive and impractical where the size of vocabulary is large.
@@
In our implementation, we use quantum states as word vectors, and use quantum fidelity to apply cosine similarity. That is, encoding two words as and , we measure their similarity as via the swap test (Barenco et al., 1997) as shown in Figure 8.
One of the challenges of quantum word embedding is how to efficiently load words as quantum states. We consider two potential schemes for embedding words to quantum state: the memory-efficient embedding and the circuit-efficient embedding. In the memory-efficient embedding, the quantum state of every word in vocabulary of size is represented by a single unitary operation , where and is a set of learnable parameters. The -qubit quantum state for the -th word is obtained from by applying it to the computational basis state and discarding ancillary qubits.
This scheme allows us to store a large number of words in small number of qubits, which is exponentially efficient in memory usage. However, the resulting quantum circuit that has sufficient expressiveness to implement has exponential depth, making it impractical for circuit-based quantum computation. Moreover, the state preparation process involves post-selection and is thus non-deterministic due to the probabilistic nature of measurement.
In contrast, the second scheme, circuit-efficient embedding, represents -th word by a quantum state of the form , where is a unitary operation parameterized by , which is specific to each word. This allow us to prepare the quantum state using a depth-efficient circuit in a deterministic process, without using excessive ancillary qubits. While it requires more classical memory to store the parameters, it is more flexible since one can add or remove words from the vocabulary during training.
To learn the parameters for these embedding schemes, we adapt the classical word2vec methods, CBOW, Skip-gram and SGNS, to the quantum setting. For quantum CBOW and Skip-gram, we introduce a parameterized unitary operator that defines the probability distribution , where is either a pooled (e.g. averaged) embedding of context words for CBOW or the embedding for the target word for Skip-gram. This removes the need of computing the costly softmax function by using the natural output of the quantum circuit to predict the index among computational basis states.
We note that quantum CBOW, cannot be directly applied for quantum word embeddings since the direct averaging of quantum states is not a natural operation for quantum computers. Instead, it may be possible to use superpositions of quantum states.
@@ 4.2 Using Quantum Embeddings
Once words are encoded as vectors, these vectors can be used in many machine learning systems, including support vector machines, which can be used for supervised classification. This sometimes involves calculating a ‘kernel’ function, which computes similarities between input vectors, sometimes involving computations that would be intractable if all the coordinates were constructed and compared explicitly (Géron, 2019, Ch 5). This is regarded as a promising research direction for quantum machine learning, because ‘quantum kernel’ circuits that compare coordinates or amplitudes can be implemented using just qubits (Schuld and Petruccione, 2021). However, as with the QPOSTR string encoding example in Section 3, if the number of gates still scales with the number of coordinates, the use of a logarithmic number of qubits is a saving of space but with no corresponding time advantage.
Several quantum vector encodings or ‘feature maps’ for word embedding vectors were compared by Alexander and Widdows (2022), and used for sentiment analysis experiments. The ZZ-feature map was found to be the most successful, achieving a classification accuracy of 62% on classification experiments involving small test sets of roughly 10K words each. This result showed initial promise, and was the largest quantum text experiment reported to-date, but also indicates how small today’s quantum NLP experiments, are compared with even modest-sized classical NLP systems.
@@ 5 Sequential Models for Text Generation
A class of Bayesian network models, called n-gram models, have been successful in multiple language processing tasks, including information retrieval, text generation, and part-of-speech tagging (Jurafsky and Martin, 2023). However, they suffer from poor performance and generalizability issues in sparse data regimes and fail to capture nonlocal syntactic relations. Handling out-of-vocabulary words or resolving ambiguity also pose challenges as n-grams do not have built-in semantic understanding (Bengio et al., 2000). A Hidden Markov models (HMM) is a Markov model whose output from any given state is probabilistic rather than deterministic, which ‘hides’ the internal state. A canonical example is when the hidden states are part-of-speech tags, such N(oun) or V(erb), which generate explicit words with given probabilities (Manning and Schütze, 1999, §9.2).
The feed forward and recurrent neural networks are specific instances of the HMM. However, HMMs also face challenges in estimating accurate probabilities when the context size increases or when the correlations get too long in languages. Beyond language processing, HMS have been widely used in other scientific fields as well. In Durbin et al. (1998), the authors discuss probabilistic models, particularly HMMs and their derivatives, and their applications in biological sequence analysis. Language models, particularly those developed for aligning and comparing sequences, can be designed to recognize patterns in biological sequences, infer evolutionary relationships, and identify functional elements.
@@ 6 Attention in Quantum NLP Models
So far we have discussed models for studying sentences as word / token sequences. Making such models scale to longer sequences has always been a challenge: with -gram models, the value of has always been small (Manning and Schütze, 1999, Ch 6); and RNN architectures including LSTM’s, while accurate for short sequences, had trouble scaling to cover long-range dependencies (Géron, 2019, Ch 15).
Attention is an additional component or technique designed to address this problem. The attention methodology was used to enhance an RNN sequence model for machine translation by Bahdanau et al. (2014), which enabled the model to capture longer-range relationships as well. Although it still relied on the encoder-decoder paradigm, the bidirectional RNN architecture introduced in Bahdanau et al. (2014) features a distinct context vector for each word in the sentence. Each context vector depends on a sequence of annotations which contains information about the entire sentence with a strong focus on the parts of the sentence surrounding the context vector’s associated input word. The annotations are weighted according to an alignment model, which scores how well an output token matches inputs around a given position.
Vaswani et al. (2017) developed this approach further, demonstrating a system where “attention is all you need.” This Transformer model — centered around scaled dot-product attention — made previous RNN-based encoder-decoder architectures obsolete when it demonstrated improved performance on various translation tasks.
@@
In particular, the “Scaled Dot-Product” attention layer featured in Vaswani et al. (2017) computes the dot-products of the query with all the keys, normalizes according to the dimension of the query and key vectors, and then applies the softmax function to obtain the weights of all the pairs, which are the values.
@@
A key advantage enjoyed by the Transformer over the previous RNN architectures is that the former can be more easily parallelized; in particular, the attention coefficients can be parallelized not only across training examples, but also over pairs of tokens in a sequence. In addition, the computational cost does not depend on the distance between tokens in the sequence, as in previous models. Together, these properties accounted for a drastic reduction in training time over sequential RNN models. The main drawback is that the computational complexity still scales quadratically in the number of tokens in a given sequence (roughly the number of words in a sentence). The problem of approximating or providing an alternative to self-attention with subquadratic complexity spawned its own burgeoning research field Zaheer et al. (2020); Poli et al. (2023); nonetheless, the Transformers remains the cornerstone of contemporary LLMs and thus will be our main focus.
6.1 Survey of quantum self-attention mechanisms
@@
In hopes of improving this scaling, and since attention layers have become so successful as key components in state-of-the-art models for NLP tasks, various quantum approaches have been suggested. For the rest of this section, we discuss different proposals for leveraging quantum computers to implement attention mechanisms.
At a high level, proposals for quantum self-attention mechanisms aim to leverage quantum hardware to gain advantage in computing the compatibility between queries and keys. Two main avenues for advantage have been identified: leveraging high-dimensional Hilbert spaces where it may be easier to separate irrelevant feature vectors, and reducing computational complexity by applying quantum gates in parallel. The former builds on results like Liu et al. (2021), which suggests that encoding feature vectors into high-dimensional spaces using quantum computers can deliver advantage in classification tasks over classical methods.
Li et al. (2022) claims to offer the first quantum self-attention implementation, QSANN. This proposal seeks quantum advantage along the first direction: by mapping encoded feature vectors into a high-dimensional Hilbert space using a quantum circuit, QSANN aims to extract correlations that are intractable classically.
@@
Concretely, Li et al. (2022) implement their self-attention mechanism as follows. For an illustration see Figure 12. First they construct an encoder circuit to load classical feature vectors onto an -qubit quantum state; they use one classical feature vector for each token in the input sequence. The number of qubits is a hyper-parameter that should be adjusted as relevant to available hardware. Next they apply parametrized quantum circuits, with identical gate layouts but different parameter values in order to compute the query, key, and value vectors for each classical feature vector. The circuit layout is illustrated in Figure 13. At this stage the query, key, and value vectors are encoded as quantum states, so measurements must be made in order to extract useful information; the resulting query and key are the expectation values of the Pauli- operators applied to the first qubit of the resulting states, and the value is a vector of expectation values of various Pauli operators. Then attention scores are then computed on a classical device as a weighted average of the output values. Interestingly, Li et al. (2022) introduce a Gaussian kernel to compute the weights on the values vector; they claim the Gaussian kernel can more easily correlate quantum states with little overlap, which is needed, e.g., if two tokens are closely related in a sentence but their quantum state embeddings happen to be distant in the qubit state space.
@@
We note that Li et al. (2022)’s proposal still requires quadratic classical computation, and its main source of quantum advantage relies on using efficiently processing vectors in high-dimensional Hilbert space to unearth hidden relationships between embedded tokens. Zhao et al. (2022) builds on these ideas, seeking quantum advantage in the same vein. By introducing various sets of ancilla qubits, the authors obviate the need to perform intermediate measurements during the attention computation. (This could be thought of as a more sophisticated example of the ancilla ‘readout qubits’ pattern used in the QPOSTR design of Section 2.) In this modality, query, key, and value quantum state-vectors are computed by applying parametrized ansatze and swapping onto ancilla registers sequentially, as shown in Figure 14. Compatibility between query and keys is computed by a Quantum Logical Similarity (QLS) module, which is implemented as a sequence of Toffoli and CNOT gates, as shown in Figure 15. This is a key step: it computes the overlap between query on keys directly on the quantum device, thereby improving on Li et al. (2022)’s result.
@@
While these two proposals address quantum self-attention mechanisms in QNLP directly, we note that Cherrat et al. (2022) proposes one for use in Vision Transformers for image classification. It’s interesting to note that the authors there seek quantum advantage in reducing the computational cost of the scaled dot-product attention calculation. Concretely, the authors introduce so-called orthogonal layers to compute compatibility scores between query and keys on the quantum hardware; these layers efficiently implement parametrized transformations on encoded feature vectors, as described in Figure 16. The main novelty here is that Cherrat et al. (2022) use the unary encoding circuit to encode token feature vectors into the Hamming weight-1 subspace of the qubit state space. This encoding is advantageous because their orthogonal layers preserve the subspace and they can be used to compute dot-products between query and keys in logarithmic time, assuming quantum gates can be applied in parallel.
@@ 7 Syntactic Parsing and Logical Forms
The use of general AI techniques such as RNNs and attention has fueled much recent success with NLP, partly because it has enabled much cross-fertilization between language and other kinds of data such as images, audio, and graphs. There are also more traditional NLP techniques based on grammatical structures found particularly in human language.
@@
Natural language (and artificial programming languages) express many structured relationships that go beyond proximity in a sequence model. For example, in the sentence “Kim kicked the ball into the goal from half-way down the pitch, right past the goalkeeper, and scored”, the grammatical structure of English enables us to infer easily that the person who scored is Kim, in spite of there being 16 words (including 4 other nouns) in between the noun Kim and the verb scored. Language grammar and syntax studies how these relationships are expressed and structured in different languages, and this field powerfully influenced much of theoretical and computational linguistics during the second-half of the century, particularly through the work of Chomsky (1957, 1965). In such a framework, syntax is the central generative system, on which other aspects of language like phonology and semantics depend (Jackendoff, 2002, Ch 5).
@@
During the century, the reliance of NLP techniques on grammatical rules has declined. This is partly due to the success of statistical and machine-learned models that share methods with other data-intensive areas of AI, and perhaps because the increasing preponderance of informal text created on smartphones immediately contradicts any assumption that the input to an NLP system should consist of distinct grammatical sentences. In computational terms, best-performing parsing algorithms have included the CYK-parser which is worst-case , while the more general attention mechanism discussed in the previous section has a baseline performance, which robustly enables more resources to be targeted at important long-range relationships. Hence the standard NLP toolbox in 2024 includes things like vectors and tensors, neural networks, convolution layers, attention heads, and gradient descent, much more typically than language syntax, grammar, phrase structure and attachment: so in most large-scale NLP systems today, a grammatical parser is not an explicit component. When grammatical parsing is discussed, it is often as a historical challenge that LLMs can solve quite effectively, not a contemporary challenge for building language models in the first place (Min et al., 2023).
@@
However, the use of grammatical parsers has been re-introduced in parts of quantum NLP, motivated especially by a mathematical correspondence between the compositional rules of tensors and categorial grammars demonstrated by Coecke et al. (2010). This framework is implemented in the lambeq system of Kartsaklis et al. (2021) which is particularly designed for quantum computers (Lorenz et al., 2021). In this system, a grammatical parser prepares a parse-tree from a sentence, which is structurally mapped to a tensor network which is compiled into a quantum circuit: so the quantum circuit encodes the grammatical dependencies of the sentence, assuming its input data is uniquely-parsed sentences.
@@
The parsing problem itself is also an interesting challenge for quantum computing, for combinatoric reasons. Computationally, the problem can be phrased as taking a list of words as input, and returning a ‘parse tree’, which is a data structure saying how the elements of the sentence are grouped together, and what grammatical role they play. The two possible tree-structures for a 3-word sentence are shown in Figure 17. For the simple case of a 3-word sentence, there are only 2 possible trees, one for the structure and one for the structure . For a 4-word sentence, the five possible trees are are shown in Figure 18. The number of possible parse trees for a sentence of length grows exponentially and is given by the Catalan number .
@@
However, a key challenge with the parsing problem is that we start with labels only for the leaf nodes, and until there is some hypothesis for the which internal nodes represent which syntactic chunks, there are no estimates for the weights on internal links. (This is part of the motivation for the use of dynamic programming in classical parsers.) Combinatorically, the problem is not exactly to find a minimum spanning tree, because internal nodes that do not correspond to distinct grammatical phrases or ‘constituents’ do not need to be visited. (Moreover, the initial graph is directed: this precludes solutions where the leaf nodes are all visited by a zigzag path along the bottom, because upward steps are forbidden.)
@@
One potential benefit of quantum parsing is that a quantum system may represent more nuanced proposals than just returning a single best parse, or even a classical mixture (estimated probability distribution over different discrete parse proposals). Rather than insisting that all language inputs must ‘pass’ a parsing test before being processed, a syntactic parser that exchanges quantum information with semantic components in parallel may become a different kind of asset in more parallel architectures.

### v2

@@ 4.1 Building Quantum Embeddings
Next we discuss techniques relating to embeddings intended for use on actual quantum devices, treating separately the topics of building quantum embeddings and using them. The circuits proposed in this section are intended for NISQ-era rather than fault-tolerant devices.
One of the most well-known recent techniques for building word embeddings is word2vec (Mikolov et al., 2013a). Taking inspiration from this line of work, we propose a quantum computing implementation of word2vec.
Word2vec is a group of word embedding methods that use shallow neural networks to capture the semantic properties of words. Word2vec includes two popular methods: Continuous Bag-of-Words (CBOW) and Skip-gram. CBOW and Skip-gram have different ways of learning the word embeddings. CBOW takes the context words around the target word as input and tries to predict the target word. Skip-gram does the opposite: it takes the target word as input and tries to predict the context words. The output of both methods is a probability distribution over all of the words in the vocabulary, which is computed using the softmax function. This can be computationally expensive and impractical when the size of vocabulary is large.
@@
In our implementation, we use quantum states as word vectors, and use quantum fidelity to apply cosine similarity. That is, encoding two words as and , we measure their similarity as via the swap test (Barenco et al., 1997), as shown in Figure 8.
One of the challenges of quantum word embedding is how to efficiently load words as quantum states. We consider two potential schemes for embedding words to quantum state: memory-efficient embedding and circuit-efficient embedding.
In memory-efficient embedding, the quantum state of every word in vocabulary of size is represented by a single unitary operation , where and is a set of learnable parameters. The -qubit quantum state for the -th word is obtained from by applying it to the computational basis state and discarding ancillary qubits.
This scheme allows us to store a large number of words in small number of qubits, which is exponentially efficient in memory usage. However, the resulting quantum circuit that has sufficient expressiveness to implement has exponential depth, making it impractical for circuit-based quantum computation. Moreover, the state preparation process involves post-selection, and is thus non-deterministic due to the probabilistic nature of measurement.
In contrast, the second scheme, circuit-efficient embedding, represents the -th word by a quantum state of the form , where is a unitary operation parameterized by , which is specific to each word. This allows us to prepare the quantum state using a depth-efficient circuit in a deterministic process, without using excessive ancillary qubits. While it requires more classical memory to store the parameters, it is more flexible since one can add or remove words from the vocabulary during training.
The circuit-efficient and memory-efficient patterns are depicted in Figures 9 and 10. In structure, the circuit-efficient pattern is like the word-embedding in word2ket (Panahi et al., 2019), and the memory-efficient pattern is like the whole vocabulary encoding. For word2ket, the motivation for expressing a whole vocabulary in a more entangled tensor network is to reduce classical memory, whereas in this design, it makes more use of the efficient quantum memory.
To learn the parameters for these embedding schemes, we adapt the classical word2vec methods, CBOW, Skip-gram and SGNS, to the quantum setting. For quantum CBOW and Skip-gram, we introduce a parameterized unitary operator that defines the probability distribution , where is either a pooled (e.g. averaged) embedding of context words for CBOW, or the embedding for the target word for Skip-gram. This removes the need of computing the costly softmax function, by using the natural output of the quantum circuit to predict the index among computational basis states. We note that quantum CBOW, cannot be directly applied for quantum word embeddings, since the direct averaging of quantum states is not a natural operation for quantum computers. Instead, it may be possible to use superpositions of quantum states.
@@ 4.2 Using Quantum Embeddings
Once words are encoded as vectors, these vectors can be used in many machine learning systems, including support vector machines, which can be used for supervised classification. This sometimes involves calculating a kernel function, which computes similarities between input vectors, sometimes involving computations that would be intractable if all the coordinates were constructed and compared explicitly (Géron, 2019, Ch 5). This is regarded as a promising research direction for quantum machine learning, because quantum kernel circuits that compare coordinates or amplitudes can be implemented using just qubits (Schuld and Petruccione, 2021). However, as with the QPOSTR string encoding example in Section 3, if the number of gates still scales with the number of coordinates, the use of a logarithmic number of qubits is a saving of space but with no corresponding time advantage.
Several quantum vector encodings or “feature maps” for word embedding vectors were compared by Alexander and Widdows (2022), and used for sentiment analysis experiments. The ZZ-feature map was found to be the most successful, achieving a classification accuracy of 62% on classification experiments involving small test sets of roughly 10K words each. This result showed initial promise, and was the largest quantum text experiment reported to-date, but also indicates how small today’s quantum NLP experiments, are compared with even modest-sized classical NLP systems.
@@ 5 Sequential Models for Text Generation
A class of Bayesian network models, called n-gram models, have been successful in multiple language processing tasks, including information retrieval, text generation, and part-of-speech tagging (Jurafsky and Martin, 2023). However, they suffer from poor performance and generalizability issues in sparse data regimes and fail to capture nonlocal syntactic relations. Handling out-of-vocabulary words or resolving ambiguity also pose challenges as n-grams do not have built-in semantic understanding (Bengio et al., 2000). A Hidden Markov models (HMM) is a Markov model whose output from any given state is probabilistic rather than deterministic, which hides the internal state. A canonical example is when the hidden states are part-of-speech tags, such N(oun) or V(erb), which generate explicit words with given probabilities (Manning and Schütze, 1999, §9.2).
The feed forward and recurrent neural networks are specific instances of the HMM. However, HMMs also face challenges in estimating accurate probabilities when the context size increases or when the correlations get too long in languages. Beyond language processing, HMMs have been widely used in other scientific fields as well. In Durbin et al. (1998), the authors discuss probabilistic models, particularly HMMs and their derivatives, and their applications in biological sequence analysis. Language models, particularly those developed for aligning and comparing sequences, can be designed to recognize patterns in biological sequences, infer evolutionary relationships, and identify functional elements.
@@ 6 Attention in Quantum NLP Models
So far we have discussed models for studying sentences as word / token sequences. Making such models scale to longer sequences has always been a challenge: with -gram models, the value of has always been small (Manning and Schütze, 1999, Ch 6); and RNN architectures including LSTMs, while accurate for short sequences, had trouble scaling to cover long-range dependencies (Géron, 2019, Ch 15).
6.1 Attention in Classical LLMs
Attention is designed to address this problem. The attention methodology was used to enhance an RNN sequence model for machine translation by Bahdanau et al. (2014), which enabled the model to capture longer-range relationships as well. Although it still relied on the encoder-decoder paradigm, the bidirectional RNN architecture introduced in Bahdanau et al. (2014) features a distinct context vector for each word in the sentence. Each context vector depends on a sequence of annotations which contains information about the entire sentence with a strong focus on the parts of the sentence surrounding the context vector’s associated input word. The annotations are weighted according to an alignment model, which scores how well an output token matches inputs around a given position.
Vaswani et al. (2017) developed this approach further, demonstrating a system where transformer blocks incorporating attention, layer norm, multi-layer perceptrons, and residual connections, fully replace recursive units. This Transformer model — centered around scaled dot-product attention — made previous RNN-based encoder-decoder architectures obsolete when it demonstrated improved performance on various translation tasks.
@@ 6.1 Attention in Classical LLMs
In particular, the “Scaled Dot-Product” attention layer featured in Vaswani et al. (2017) computes the dot-products of the query with all the keys, normalizes according to the dimension of the query and key vectors, and then applies the softmax function to obtain the weights of all the pairs, which are the values. Rather then enumerate all the indices for summation, it is typical to write the lists of vectors as matrices, whereby the definition takes the common form
where is the embedding dimension, , , and are matrices of size where is the number of words / tokens in a sequence, and the softmax function is applied to each row. This formulates dot product attention as a matrix multiplication (), a softmax step (), and a final matrix multiplication ().
@@
A key advantage enjoyed by the Transformer over the previous RNN architectures is that this multiplication can be parallelized, which computes the pairwise relationships between all the tokens in a sequence at once. In addition, the computational cost does not depend on the distance between tokens in the sequence, as in previous models. Together, these properties accounted for a drastic reduction in training time over sequential RNN models. The main drawback is that the computational complexity still scales quadratically in the number of tokens in a given sequence (roughly the number of words in a sentence). The problem of approximating or providing an alternative to self-attention with subquadratic complexity spawned its own burgeoning research field (Zaheer et al., 2020; Poli et al., 2023).
6.2 Near-term quantum self-attention mechanisms
@@
In hopes of improving this quadratic scaling, and since attention layers have become so successful as key components in state-of-the-art models for NLP tasks, various quantum approaches have been suggested. This section focuses on near-term quantum circuit designs.
@@
Li et al. (2022) claim to offer the first quantum self-attention implementation, QSANN. By mapping encoded feature vectors into a high-dimensional Hilbert space using a quantum circuit, QSANN aims to extract correlations that are intractable classically. For an illustration see Figure 14. First they construct an encoder circuit to load classical feature vectors onto an -qubit quantum state; they use one classical feature vector for each token in the input sequence. The number of qubits is a hyper-parameter that should be adjusted as relevant to available hardware. Next they apply parametrized quantum circuits, with identical gate layouts but different parameter values in order to compute the query, key, and value vectors for each classical feature vector. The circuit layout is illustrated in Figure 15. At this stage the query, key, and value vectors are encoded as quantum states, so measurements must be made in order to extract useful information; the resulting query and key are the expectation values of the Pauli- operators applied to the first qubit of the resulting states, and the value is a vector of expectation values of various Pauli operators. Attention scores are then computed on a classical device as a weighted average of the output values. Interestingly, Li et al. (2022) introduce a Gaussian kernel to compute the weights on the values vector; they claim the Gaussian kernel can more easily correlate quantum states with little overlap, which is needed, e.g., if two tokens are closely related in a sentence but their quantum state embeddings happen to be distant in the qubit state space. The proposal of Li et al. (2022) still requires quadratic classical computation, and its main source of quantum advantage relies on using efficiently processing vectors in high-dimensional Hilbert space to unearth hidden relationships between embedded tokens.
@@
The work of Zhao et al. (2022) builds on these ideas, seeking quantum advantage in the same vein. By introducing various sets of ancilla qubits, the authors obviate the need to perform intermediate measurements during the attention computation. (This could be thought of as a more sophisticated example of the ancilla readout qubits pattern used in the QPOSTR design of Section 2.) In this modality, query, key, and value quantum state-vectors are computed by applying parametrized ansatze and swapping onto ancilla registers sequentially, as shown in Figure 16. Compatibility between query and keys is computed by a Quantum Logical Similarity (QLS) module, which is implemented as a sequence of Toffoli and CNOT gates, as shown in Figure 17. This is a key step: it computes the overlap between query on keys directly on the quantum device, thereby improving on Li et al. (2022).
@@
While these two proposals address quantum self-attention mechanisms in QNLP directly, Cherrat et al. (2022) proposes one for use in Vision Transformers for image classification, seeking quantum advantage in reducing the computational cost of the scaled dot-product attention calculation. Concretely, the authors introduce so-called orthogonal layers to compute compatibility scores between query and keys on the quantum hardware; these layers efficiently implement parametrized transformations on encoded feature vectors, as described in Figure 18. The main novelty here is that Cherrat et al. (2022) use the unary encoding circuit to encode token feature vectors into the Hamming weight-1 subspace of the qubit state space. This encoding is advantageous because their orthogonal layers preserve the subspace, and they can be used to compute dot-products between query and keys in logarithmic time, assuming quantum gates can be applied in parallel. Cherrat et al. (2022) report preliminary results from simulation, and with a 6-qubit quantum processor.
6.3 Attention on Fault-Tolerant Quantum Computers
The role of attention in systems like GPT (Brown et al., 2020) has spurred more ambitious proposals, including the recent preprints of Liao and Ferrie (2024) and Guo et al. (2024), which describe large-scale versions of full transformer-inspired processes on fault-tolerant error-correcting quantum computers. Guo et al. (2024) use quantum signal processing and singular value transformation techniques to apply a polynomial approximate softmax, whereas (Liao and Ferrie, 2024), like (Zhao et al., 2022), replace softmax with a linear quantum alternative.
A main challenge for such proposals remains quantum I/O, which these papers acknowledge, but do not solve explicitly. To extract all the weights from a matrix stored in a quantum state would typically require many more shots than there are weights. A smaller alternative is to output token indices directly. A natural NISQ-inspired suggestion is to adapt the token index generation approaches from Section 5 to transformer prototypes. While adding a token-generation head may be useful for circumventing I/O issues, the complexity and depth of the remaining bodies of the circuits in these proposals puts practical implementations beyond the NISQ era.
A quantum decoder, for optimized search through a much larger space of long proposed token sequences, is proposed by Bausch et al. (2021), which casts the problem as probabilistically branching tree-search, which is mathematically equivalent to probabilistic grammar parsing. Even though the hardware capabilities for such quantum operations are still some years away, this connects optimizations in sequence generation from today’s softmaxed probabilities, with traditional syntactic approaches to language modeling, which we discuss next.
@@ 7 Syntactic Parsing and Logical Forms
The use of general AI techniques such as RNNs and attention has fueled much recent success with NLP, partly because it has enabled much cross-fertilization between language and other kinds of data such as images, audio, and graphs. There are also more traditional NLP techniques, based on grammatical structures found particularly in human language.
@@
Natural languages (and artificial programming languages) express many structured relationships that go beyond proximity in a sequence model. For example, in the sentence “Kim kicked the ball into the goal from half-way down the pitch, right past the goalkeeper, and scored”, the grammatical structure of English enables us to infer easily that the person who scored is Kim, in spite of there being 16 words (including 4 other nouns) in between the noun Kim and the verb scored. Language grammar and syntax studies how these relationships are expressed and structured in different languages, and this field powerfully influenced much of theoretical and computational linguistics during the second-half of the century, particularly through the work of Chomsky (1957, 1965). In such a framework, syntax is the central generative system, on which other aspects of language like phonology and semantics depend (Jackendoff, 2002, Ch 5).
@@
During the century, the reliance of NLP techniques on grammatical rules has declined. This is partly due to the success of statistical and machine-learned models that share methods with other data-intensive areas of AI, and perhaps because the increasing preponderance of informal text created on smartphones immediately contradicts any assumption that the input to an NLP system should consist of distinct grammatical sentences. In computational terms, best-performing parsing algorithms have included the CYK-parser which is worst-case , while the more general attention mechanism discussed in the previous section has a baseline performance, which robustly enables more resources to be targeted at important long-range relationships. Hence in most large-scale NLP systems today, a grammatical parser is not an explicit component: and with the challenge of reducing computational costs below quadratic, there would need to be a strong reason for requiring the extra burden of a cubic computing step for any large-scale model. When grammatical parsing is discussed, it is often as a historical challenge that LLMs can solve quite effectively, not a contemporary challenge for building language models in the first place (Min et al., 2023).
@@
However, the use of grammatical parsers has been re-introduced in parts of quantum NLP, motivated especially by a mathematical correspondence between the compositional rules of tensors and categorial grammars demonstrated by Coecke et al. (2010). This framework is implemented in the lambeq system of Kartsaklis et al. (2021) which is particularly designed for quantum computers (Lorenz et al., 2023). In this system, a grammatical parser prepares a parse-tree from a sentence, which is structurally mapped to a tensor network which is compiled into a quantum circuit: so the quantum circuit encodes the grammatical dependencies of the sentence, assuming its input data consists of uniquely-parsed sentences.
@@
The parsing problem itself is also an interesting challenge for quantum computing, for combinatoric reasons. Computationally, the problem can be phrased as taking a list of words as input, and returning a parse-tree, which is a data structure saying how the elements of the sentence are grouped together, and what grammatical role they play. The two possible tree-structures for a 3-word sentence are shown in Figure 19. For the simple case of a 3-word sentence, there are only 2 possible trees, one for the structure and one for the structure . For a 4-word sentence, the five possible trees are are shown in Figure 20. The number of possible parse trees for a sentence of length grows exponentially and is given by the Catalan number .
@@
However, a key challenge with the parsing problem is that we start with labels only for the leaf nodes, and until there is some hypothesis for the which internal nodes represent which syntactic chunks, there are no estimates for the weights on internal links. (This is part of the motivation for the use of dynamic programming in classical parsers.) Combinatorically, the problem is not exactly to find a minimum spanning tree, because internal nodes that do not correspond to distinct grammatical phrases or constituents do not need to be visited. (Moreover, the initial graph is directed: this precludes solutions where the leaf nodes are all visited by a zigzag path along the bottom, because upward steps are forbidden.)
@@
One potential benefit of quantum parsing is that a quantum system may represent more nuanced proposals than just returning a single best parse, or even a classical mixture (estimated probability distribution over different discrete parse proposals). Rather than insisting that all language inputs must pass a parsing test before being processed, a syntactic parser that exchanges quantum information with semantic components in parallel may become a different kind of asset in more parallel architectures.